OpenStudy (anonymous):

whats the solution set for |2x-6|-7=7

OpenStudy (anonymous):

To find the solution set, solve two equations: 2x-6-7=7 2x-13=7 2x=21 x=21/2 and -(2x-6)-7=7 -2x+6-7=7 -2x-1=7 -2x=8 x=-4 So, the solution set is [-4, 21/2]